OpenStudy (beccab003):

MEDAL TO BEST ANSWER Given three consecutive integers with a sum of four times the second number, write an algebraic equation for this situation. n + n + 1 +n + 2 = 4n n + n + 1 + n + 2= = 4(n+1) 3n+3=4n

OpenStudy (beccab003):

@Directrix. Last question if you could help.

Directrix (directrix):

What is all this, your answer? n + n + 1 +n + 2 = 4n n + n + 1 + n + 2= = 4(n+1) 3n+3=4n

OpenStudy (beccab003):

It is the answer options

OpenStudy (beccab003):

okay. I thought it was the last one, but, based on a different problem, that I can post if you want, the answer was similar to the second one.

Directrix (directrix):

I had to finish up a problem I had helped a guy get started on before you called me.

Directrix (directrix):

The three consecutive integers are: n n + 1 n + 2

OpenStudy (beccab003):

Here is the problem that is similar that also has my teacher's feedback. Given five consecutive integers with a sum of three times the second number, write an algebraic equation for this situation. 5n + 5 = 3n ----answer I chose n + n + 1 + n + 2 + n + 3 + n + 4 = 3(n + 1) --correct answer (n + 1)^5 = 3n N + n + 1 + n + 2 + n + 3 + n + 4 = 3n FEEDBACK: When you translate this expression into math terms, the result is: n + n + 1 + n + 2 + n + 3 + n + 4 = 3(n + 1) Rebember that, given that the first term is n, the consecutive integers are ( n + 1), (n + 2), and so on. Therefore, the “second number” is (n+1)

OpenStudy (beccab003):

Hopefully that helps. I think it is the second answer but I need to be sure.

Directrix (directrix):

Now, we have to translate the words of the problem to mathematical symbols. three consecutive integers with a sum ====================== n + n+1 + n + 2 a sum of four times the second number, ========================= 4*(n+1)

Directrix (directrix):

How (mathematical operation) are these two expressions related? n + n+1 + n + 2 and 4*(n+1)

Directrix (directrix):

@BeccaB003

OpenStudy (beccab003):

They are related because it is a sum of four numbers times the second number which is n...

Directrix (directrix):

So, what equation would you write to illustrate that?

OpenStudy (beccab003):

n + n + 1 + n + 2= 4(n+1)

Directrix (directrix):

Correct.

OpenStudy (beccab003):

thanks!

Directrix (directrix):

You're welcome.